OpenStudy (anonymous):

If cos(x)=-5/13, and x satisfies 0

OpenStudy (anonymous):

i think you just did this exact problem \[\cos(x)=-\frac{5}{13}\] \[\sin(x)=\frac{12}{13}\] and now double angle formulas

OpenStudy (jamesj):

i.e., cos 2x = cos^2 x - sin^2 x sin 2x = 2 sin x . cos x ...